OpenStudy (anonymous):

Can you please explain and solve this problem for Q?: 7.8 x 10^6 = [(9.0 x 10^9) x (Q) x (2Q)]/0.020^2 Can you please explain and solve this problem for Q?: 7.8 x 10^6 = [(9.0 x 10^9) x (Q) x (2Q)]/0.020^2 @Physics

OpenStudy (anonymous):

\[ 7.8 \times 10^6 = [(9.0 \times 10^9) \times (Q) \times (2Q)]/0.020^2\]Re-arrange the equation by first multiplying both sides by \[0.020^2\]then dividing both sides by \[9.0 \times10^9\] and then dividing both sides by \[2\] to give \[ (7.8 \times 10^6) \times (0.020^2)/(9.0 \times 10^9)/2 = [ (Q) \times (Q)]\]Lets look at the right hand side of the equation now which is \[ (Q) \times (Q)]\]which simplifies to \[Q^2\]so square root both sides of the equation to give\[Q=\sqrt{(7.8 \times 10^6) \times (0.020^2)/(9.0 \times 10^9)/2}\] which is \[Q=4.16 \times 10^{-4}\] to 3 significant figures. Remember to include your units where necessary.

OpenStudy (anonymous):

aha! thank you SOO much! that helped a ton:D